# deepSURF: Detecting Memory Safety Vulnerabilities in Rust Through Fuzzing LLM-Augmented Harnesses

### v1

@@ 4.2 Arguments Analysis
Rust supports a diverse range of types, from primitive types such as integers to common types found in other programming languages, such as structs and generics, as well as Rust-specific types like closures. deepSURF is designed to handle this variety to enable effective fuzzing of URAPIs with diverse inputs, addressing the limitations of state-of-the-art tools in supporting different argument types. A complete list of the data types that deepSURF can generate using fuzzer’s input is provided in Appendix A.
@@ 4.2.3 Recursive Type Analysis
To invoke URAPIs, the harness must construct argument types from the fuzzer’s byte stream input. While trivial for primitives, this is more complex for structs, generic types, and container types (e.g., vectors). To ensure correct invocation of library functions, deepSURF performs recursive type analysis on function arguments during static analysis and leverages rustc’s type-checking system for precise type matching. For generic and complex types, deepSURF analyzes trait function or constructor arguments, respectively. For container types, it recursively analyzes inner types. The pseudocode for this recursive analysis is provided in Appendix A.
@@
3
4impl STrait for CustomTy0 {
@@
6 if _to_u8(fz_data)%2 == 0{ panic!("INTENTIONAL PANIC!") }
7 return _to_string(fz_data);
@@
11
12fn main (){
13 fuzz!(|fz_data: &[u8]| {
@@
15 let t1 = _to_string(fz_data);
16 let t2 = CustomTy0(t1);
17 /* Similar steps with the above for building t3*/
18 &t0.foo(t2, t3);
@@ 4.3 Static Harness Generation
Generating Arguments. For primitive type nodes, deepSURF directly converts the fuzzer’s input bytes into the required types using helper functions. For generic nodes, it substitutes all occurrences of the generic parameter with custom concrete types. For complex types, it first generates the constructor arguments, which uses to invoke the constructors. Listing 2(c) shows a harness generated for the foo function: the complex type Shape is instantiated via its zero constructor (line 14), while the generic parameters S and U are replaced with CustomTy0 and CustomTy1, respectively—custom structs that enable trait implementation.
@@ 4.4 LLM-based Augmentation
The harnesses generated in the previous stage represent a foundation for targeting the respective URAPIs, but they do not fully address all the challenges we explained in Section I. Specifically, they lack support for initializing complex types using multiple constructors (C2), fail to substitute generic parameters with unsafe trait bounds using library-defined concrete types (C3), and do not incorporate semantically related API call sequences that involve the targeted URAPI (C5). These limitations are addressed in the current stage using LLMs.
Figure 3 outlines the steps deepSURF performs in this stage. In summary, the Harness Augmenter attempts to use an LLM to improve statically-generated harnesses, while the Harness Selector optimally combines statically-generated and LLM-generated harnesses.
Each prompt includes: (a) the fetched harness, (b) the targeted URAPI, (c) metadata from static analysis, (d) relevant documentation, and (e) augmentation instructions. The static metadata includes available constructors for complex types and the list of URAPIs defined in the library. Documentation—extracted via rustdoc—is included when input token limits allow. The instructions guide the model on how to augment the harness and outline side effects to avoid.
@@
Harnesses Selection for LLM Augmentation and Fuzzing Corpus Generation. deepSURF’s goal is to create a corpus of fuzzing harnesses that targets unsafe code (C1) and captures diverse and complex behaviors within the fuzzed crate. However, generating too many harnesses risks dispersing the fuzzer’s efforts and reducing its effectiveness. Furthermore, statically-generated and LLM-generated harnesses often reach different code paths in the fuzzed code; thus, an appropriate combination of LLM-generated and statically-generated harnesses should be included in the final Fuzzing Corpus. For these reasons, we apply specific harness selection policies both in the Harness Augmenter and in the Harness Selector to choose which Statically Generated Harnesses to augment and which harnesses to include in the final Fuzzing Corpus.
@@
Appendix B and Appendix C evaluate alternative harness augmentation and harness selection policies, and they provide further justification for the chosen policies.
@@
deepSURF dynamically tests the generated harnesses using AFL++, delegating the task of finding bug-triggering conditions to the fuzzer. While effective at detecting memory bugs in Rust, fuzzing can produce false positives from panics or asserts inserted by developers to enforce invariants or handle errors [39, 8]. To prevent such cases from being treated as crashes, deepSURF modifies afl.rs and enables ASan [38] to detect non-crashing memory violations. Its output is then used to filter out false positives caused by large memory allocations. These measures ensure that fuzzing focuses exclusively on memory corruption vulnerabilities, fully addressing C1.
@@ 5 Evaluation
In this section, we evaluate deepSURF by addressing the following three research questions:
@@ 5.1 Experimental Setup
To evaluate deepSURF’s effectiveness in detecting memory bugs and to compare it against state-of-the-art approaches, we use the ERASAN dataset. ERASAN is a Rust-specific sanitizer, and its dataset includes 27222The xcb crate is excluded due to its dependency on external software. library crates from the RustSec repository [46], each containing real-world memory safety issues [37]. We chose this dataset because it is sufficiently large, includes well-documented memory corruption bugs across diverse libraries, and has been used in prior work [30], making it a heterogeneous and unbiased foundation for deepSURF’s evaluation.
@@
We ran our experiments on a machine equipped with AMD EPYC 7B13 CPUs (112 cores, 2.2GHz) and 224GB of memory, running Ubuntu 24.04. In all experiments, we used DeepSeek-R1 with its default settings (temperature and top-p = 1.0) and set the maximum number of prompt retry attempts to 6. For static harness generation, we considered up to 4 constructors per complex type, we selected dependency trees using seeded random sampling and attempted to compile up to 50.
@@
string-intern 0 100% (2) 33.3% (8) 33.3% (1) 33.3% (1)
http 0 86.4% (109) 48% (178)
@@
Total 26 (6) 87.3% (1080) 21.8%(653) 4% (129) 3% (72)
TABLE II: Bug-finding capability of deepSURF and comparison of URAPI Coverage and number of compilable generated harnesses with other Rust fuzzing tools. — signifies cases where the tool crashes and ! denotes cases where harness generation did not finish within 24 hours.
5.2 Evaluation Results
RQ1: deepSURF’s Bug-Finding Capability. In this experiment, we evaluate deepSURF on ERASAN’s dataset to determine its ability to automatically generate harnesses that uncover existing or new memory safety bugs. We fuzz each of the generated harnesses for 24 hours using two AFL++ threads working together on the same target: one with ASan and the other with CmpLog [20], as suggested by [1].
deepSURF generates harnesses that detect 26 memory bugs by fuzzing, as detailed in the second column of Table II. These include double-free (DF), buffer overflow (BOF), arbitrary memory accesses (SEGV), and use-after-free (UAF) violations. Out of the 26 found memory corruption vulnerabilities, 20 are previously-known which deepSURF detects automatically without any human involvement. In addition, deepSURF discovers 6 new memory bugs in three crates.
We have disclosed all the newly-discovered bugs to their affected developers. After our reporting, the bug in toodee has been patched and is pending RustSec ID assignment and the bugs in the maintained fork [^3] The crate slice-deque has been unmaintained since 2020, but the bugs we found also affect its maintained fork, slice-ring-buffer. of slice-deque are pending patches.
Many of the memory bugs detected by deepSURF involve intricate scenarios and advanced Rust features that pose challenges for automated tools. For example, triggering 12 bugs requires custom function implementations that panic or return unexpected values. Others, such as those in slice-deque and simple-slab, are triggered only by specific API call sequences. These bug-triggering patterns are typically found only in human-written PoCs, yet deepSURF generated harnesses containing them automatically.
deepSURF misses 11 reported memory bugs across the remaining 11 crates. These require language features beyond its current support, including async programming, multithreading, and sequences of function calls that involve not only library-defined APIs but also other internal operations like unexpected calls to std::mem::forget.
RQ1: deepSURF identified 26 memory safety bugs (including 6 previously-unknown) by fuzzing the harnesses that it automatically generated.
RQ2: Comparison with State-of-the-Art Tools. We evaluated existing Rust fuzzing tools on our dataset to compare their bug-finding capabilities with deepSURF. Specifically, we ran RUG [11], RPG [50] and RULF [23] to generate harnesses for the libraries in our dataset and fuzzed these targets to compare their findings with deepSURF. For this comparison, the other tools were tested with their original settings. We could not test RuMono [54] and FRIES [51], as their code was not publicly available at the time of writing.
None of the other tools were able to automatically expose any of the bugs that deepSURF uncovered in our dataset. Moreover, as shown in Table II, deepSURF achieved a high overall URAPI Coverage (see Section 2) of 87.3%, while RUG, RPG, and RULF achieved just 21.8%, 4%, and 3% respectively. This significant difference highlights deepSURF’s focus on fuzzing unsafe code reachable through APIs, along with its broad support for Rust data types, enabling it to cover a substantially larger set of potentially vulnerable APIs.
RUG was the best-performing tool among the existing ones, generating harnesses of relatively high quality. While it failed to detect any memory corruption vulnerabilities when tested with its default fuzzing engine (libFuzzer [26]), we observed that it could potentially detect one of the bugs in the toodee crate if used with AFL++. In contrast, RULF and RPG exhibit limited argument-type support, preventing them from fully addressing C2 and C3. As a result, their ability to fuzz complex libraries with diverse API argument types is limited.
None of the existing tools fully addresses C4 (support for custom implementations simulating user-defined logic) or C5 (support for generating semantically-related complex API sequences), which are critical for detecting complex memory corruption vulnerabilities. Moreover, all three tools failed to properly address C1, since they produced false positives due to crashes unrelated to actual memory safety bugs either by directly invoking unsafe code without preserving safety invariants or due to incorrect handling of panics and assertions. Finally, RUG and RPG failed to run on five crates, and RULF on one, due to crashes, timeouts, or errors caused by unimplemented features or implementation errors.

### v2

@@ 4.1.2 Detecting Public Entry Points to UEFs
TABLE II: Rust data types supported by deepSURF. N refers to an integer; T, T1, …, Tn refer to types from categories 1 to 8; Tr, Tr1, …, Trn refer to traits defining required behaviors; and E represents an error type.
Category Examples
1. Primitive Types Integers, Booleans, Floating-point numbers, Characters
2. Standard Library Types Strings, Vectors (Vec<T>), Boxes (Box<T>), Options (Option<T>), Results (Result<T, E>)
3. Slices Common (&[T]) and string slices (&str)
4. Compound Types Arrays ([T; N]), Tuples ((T1, …, Tn))
5. Complex Types Structs, Enums
7. Generic Types Type parameters (e.g., G) serve as placeholders for compatible concrete types and may require traits (G: Tr1 + … + Trn)
8. Rust-Specific Types Closures, Trait-associated Types, Dynamic Trait Objects (dyn Tr)
@@ 4.2 Arguments Analysis
Rust supports a diverse range of types, from primitive types such as integers to common types found in other programming languages, such as structs and generics, as well as Rust-specific types like trait types. deepSURF is designed to handle this variety to enable effective fuzzing of URAPIs with diverse inputs, addressing the limitations of state-of-the-art tools in supporting different argument types. A complete list of the data types that deepSURF can generate using fuzzer’s input is provided in Table II.
@@ 4.2.2 Generic Types Support
Algorithm 1 Recursive Type Analysis
1:function Analyze_Args(fn)
2:  for all $arg\in fn.args$ do
3:   $\textsc{Analyze\_Arg}(arg)$
4:  end for
5:end function
6:function Analyze_Arg(arg)
7:  $atype\leftarrow\textsc{Get\_Arg\_Type}(arg)$
8:  if $\textsc{Is\_Generic}(atype)$ then
9:   for all $fn\in\textsc{Get\_Trait\_Fns}(arg)$ do
10:    $\textsc{Mark}(fn)$; $\textsc{Analyze\_Args}(fn)$
11:   end for
12:  end if
13:  if $\textsc{Is\_Complex}(atype)$ then
14:   for all $ctor\in\textsc{Get\_Constructors}(arg)$ do
15:    $\textsc{Mark}(ctor)$; $\textsc{Analyze\_Args}(ctor)$
16:   end for
17:  end if
18:  if $\textsc{Is\_Container}(atype)$ then
19:   $\textsc{Analyze\_Arg}(\textsc{Get\_Inner\_Arg}(arg))$
20:  end if
21:  return atypeatype
22:end function
@@ 4.2.3 Recursive Type Analysis
To invoke URAPIs, the harness must construct argument types from the fuzzer’s byte stream input. While trivial for primitives, this is more complex for structs, generic types, and container types (e.g., vectors). To ensure correct invocation of library functions, deepSURF performs recursive type analysis on function arguments during static analysis and leverages rustc’s type-checking system for precise type matching. For generic and complex types, deepSURF analyzes trait function or constructor arguments, respectively. For container types, it recursively analyzes inner types. The pseudocode for this recursive analysis is provided in Algorithm 1.
4.3 Static Harness Generation
@@
3impl STrait for CustomTy0 {
@@
5 if _to_u8(DATA, 0)%2 == 0{ panic!("INTENTIONAL PANIC!") }
6 let mut t2 = _to_u8(DATA, 1) % 17;
7 return String::from(_to_str(DATA, 2, 2 + t2 as usize));
@@
11fn main (){
12 fuzz_nohook!(|data: &[u8]| {
13 if data.len() < 53 {return;}
14 set_global_data(data); // DATA = data
@@
16 let mut t7 = _to_u8(DATA, 36) % 17;
17 let t8 = _to_str(DATA, 37, 37 + t7 as usize);
18 let t9 = CustomTy0(String::from(t8));
19 /* Similar steps with the above for building t13*/
20 &t0.foo(t9, t13);
@@
Generating Arguments. For primitive type nodes, deepSURF directly converts the fuzzer’s input bytes into the required types using helper functions. For generic nodes, it substitutes all occurrences of the generic parameter with custom concrete types. For complex types, it first generates the constructor arguments, which uses to invoke the constructors. Listing 2(c) shows a harness generated for the foo function: the complex type Shape is instantiated via its zero constructor (line 15), while the generic parameters S and U are replaced with CustomTy0 and CustomTy1, respectively—custom structs that enable trait implementation.
@@ 4.4 LLM-based Augmentation
The harnesses generated in the previous stage represent a foundation for targeting the respective URAPIs, but they do not fully address all the challenges we explained in Section 3. Specifically, they lack support for initializing complex types using multiple constructors (C2), fail to substitute generic parameters with unsafe trait bounds using library-defined concrete types (C3), and do not incorporate semantically related API call sequences that involve the targeted URAPI (C5). These limitations are addressed in the current stage using LLMs.
Figure 3 outlines the steps deepSURF performs in this stage. In summary, the Harness Augmenter attempts to use an LLM to improve statically generated harnesses, while the Harness Selector optimally combines statically generated and LLM-generated harnesses.
Each prompt includes: (a) the fetched harness, (b) the targeted URAPI, (c) metadata from static analysis, (d) relevant documentation, and (e) augmentation instructions. The static metadata includes available constructors for complex types and the list of URAPIs identified in the library. Documentation—extracted via rustdoc—is included when input token limits allow. The instructions guide the model on how to augment the harness and outline side effects to avoid.
@@
Harnesses Selection for LLM Augmentation and Fuzzing Corpus Generation. deepSURF’s goal is to create a corpus of fuzzing harnesses that targets unsafe code (C1) and captures diverse and complex behaviors within the fuzzed crate. However, generating too many harnesses risks dispersing the fuzzer’s efforts and reducing its effectiveness. Furthermore, statically generated and LLM-generated harnesses often reach different code paths in the fuzzed code; thus, an appropriate combination of LLM-generated and statically generated harnesses should be included in the final Fuzzing Corpus. For these reasons, we apply specific harness selection policies both in the Harness Augmenter and in the Harness Selector to choose which statically generated harnesses to augment and which harnesses to include in the final Fuzzing Corpus.
@@
In Section 6 and Appendix A we evaluate alternative harness augmentation and harness selection policies, and we provide further justification for the chosen policies.
@@
deepSURF dynamically tests the generated harnesses using AFL++, delegating the task of finding bug-triggering conditions to the fuzzer. While effective at detecting memory bugs in Rust, fuzzing can produce false positives—either from panics or assertion failures intentionally inserted by library authors to enforce invariants and prevent incorrect API usage, or from errors in the harness itself (e.g., improper type conversions or incorrect use of unsafe code) [47, 9]. To prevent such cases from being treated as crashes, deepSURF modifies afl.rs by removing hooks that translate all panics to crashes and uses the fuzz_nohook! macro as the fuzzer’s entry point. We acknowledge that this may introduce false negatives—e.g., missed logical bugs that do not lead to memory corruption—but we accept this trade-off to reduce the manual triage required to analyze our tool’s findings. We empirically evaluate this trade-off in Appendix B.
Additionally, ASan[45] is enabled to detect non-crashing memory violations, and its output is used to automatically detect memory corruptions and exclude crashes caused by large memory allocations. These measures ensure that fuzzing focuses exclusively on memory corruption vulnerabilities, fully addressing C1.
5 Implementation
We implemented deepSURF in Rust, totaling approximately 13.1k lines of code. Static analysis extends the rustc 1.81.0-dev compiler, leveraging its query system and metadata while avoiding limitations of tools like rustdoc [51, 69]. Using a recent compiler ensures compatibility with modern libraries and fuzzing plugins [7]. Below, we detail key implementation steps of our tool.
CFG Construction and Callee Resolution. To perform unsafe reachability analysis, we construct a control flow graph (CFG) of all relevant library functions by emitting their MIR and iterating through basic blocks. Within these blocks, we identify terminators, which represent the exit points of a basic block [51]. We specifically focus on terminators corresponding to exits caused by calls to functions. By reusing the existing MIR inliner logic, we resolve these calls, classify their types (see §4.1.2), and identify locations that allow user-defined functions.
Name Resolution and Visibility. The Rust compiler assigns unique identifiers to library items like functions, traits, and types during HIR analysis [51]. However, these identifiers are internal to rustc and cannot be accessed in user code. deepSURF’s harnesses require access to these items to test the library, so we resolve these identifiers into publicly accessible names. Rust’s privacy rules and re-exports complicate this process. deepSURF overcomes these challenges by using metadata from rustc’s privacy pass [51] to derive absolute paths for the required items.
Data Type Analysis. To analyze and distinguish between data types in a Rust library, deepSURF uses ty::Ty, an internal rustc type [51]. Its precise type information enables deepSURF to perform essential tasks, such as identifying candidate constructors and analyzing the arguments of closure types.
Trait Information Gathering. Performing static analysis within rustc proved essential for the step of trait information collection. Specifically, by invoking rustc’s internal APIs, such as predicates_of and associated_items, deepSURF precisely gathers all the traits a generic parameter must implement, along with the trait items, including associated types and trait functions.
Ownership Handling. deepSURF avoids lifetime and ownership violations through specific design choices. During static harness generation, we use a “main” function as the entry point for the fuzzer, in which all URAPI arguments are local variables with lifetimes scoped to main. To preserve ownership, we avoid reusing moved values by constructing each argument from distinct segments of the fuzzer’s input. Also, we ensure that LLM-augmented harnesses pass compiler checks, including lifetime validation.
LLM Integration. deepSURF integrates DeepSeek-R1 [17], released in January 2025, due to its strong reasoning capabilities, cost-effectiveness, and open-source availability, which enables reproducibility of results. We specifically use version R1 (671 billion parameters, 164K-token context window), which offers a practical balance between performance, cost, and accessibility for research. If desired, deepSURF is configurable to use different LLMs. Moreover, library documentation is provided to the model using rustdoc-md [46].
Memory Safety of the Generated Harnesses. All harnesses generated by deepSURF and included in the Fuzzing Corpus contain only safe code, enforced by the #![forbid(unsafe_code)] directive at the beginning of each harness. This prevents false positives during fuzzing that can arise from direct calls to UAPIs that break their safety contracts, as well as including unsafe code in the harness itself. Furthermore, we verify that the use of global data complies with Rust’s safety guarantees.
@@ 6 Evaluation
In this section, we evaluate deepSURF by addressing three research questions and presenting two case studies of real-world memory corruption bugs detected by our tool. The research questions are as follows:
@@ 6.1 Experimental Setup
To evaluate deepSURF’s effectiveness in detecting memory bugs and to compare it against state-of-the-art approaches, we construct a dataset, comprising 63 crates drawn from ERAsan [35], RustSan [13], RUG [12], and CrabTree [48] datasets. ERAsan and RustSan are Rust-specific sanitizers whose datasets include crates with known real-world memory safety issues. From these datasets, we select 44 crates, excluding those that (a) for which we identified no URAPIs, (b) are duplicated across the two datasets, (c) depend on external software, or (d) would incur prohibitive LLM-generation cost. The remaining 19 crates are sourced from the datasets of the Rust testing tools CrabTree and RUG, which include popular, widely downloaded crates from crates.io. Constructing our dataset using a sufficiently large number of crates used in prior work—including many with well-documented memory bugs across diverse libraries—provides a heterogeneous and unbiased foundation for evaluating deepSURF.
@@
We ran our experiments on a machine equipped with AMD EPYC 7B13 CPUs (112 cores, 2.2GHz) and 224GB of memory, running Ubuntu 24.04. In all experiments, the default deepSURF configuration uses DeepSeek-R1 with its default settings (temperature and top-p = 1.0) and sets the maximum number of prompt retry attempts to six. For static harness generation, we considered up to four constructors per complex type, we selected dependency trees using seeded random sampling and attempted to compile up to 50. We evaluate the use of different LLMs in RQ3.
TABLE III: New memory safety bugs found by deepSURF. The first column lists the RustSec ID when available. The remaining columns report the affected crate, the vulnerable function or trait, and the bug type.
RUSTSEC ID Crate Vulnerable Function/Trait Bug Type
RUSTSEC-2025-0062 toodee toodee::DrainCol::drop BOF
RUSTSEC-2025-0044 slice_deque SliceDeque::insert DF
RUSTSEC-2025-0044 slice_deque slice_deque::IntoIter::clone DF
RUSTSEC-2025-0044 slice_deque SliceDeque::extend_from_slice DF
RUSTSEC-2025-0044 slice_deque SliceDeque::shrink_to_fit DF
- ordnung CompactVec::remove UAF
RUSTSEC-2025-0054 array-queue ArrayQueue::push_front MEMCRP
RUSTSEC-2025-0053 arenavec arenavec::common::AllocHandle MEMCRP
RUSTSEC-2025-0053 arenavec arenavec::common::allocate_inner BOF
RUSTSEC-2025-0053 arenavec arenavec::common::SliceVec::split_off DF
RUSTSEC-2025-0049 scratchpad scratchpad::Tracking BOF
RUSTSEC-2025-0050 id-map IdMap::from_iter MEMCRP
6.2 Evaluation Results
RQ1: deepSURF’s Bug-Finding Capability. In this experiment, we evaluate deepSURF on our dataset to determine its ability to automatically generate harnesses that uncover existing or new memory safety bugs. We fuzz each of the generated harnesses for 24 hours using two AFL++ threads working together on the same target: one with ASan and the other with CmpLog [22], as suggested by [1].
deepSURF generates harnesses that detect 42 memory bugs by fuzzing (see [40] for the full bug list). These include double-free (DF), buffer overflow (BOF), use-after-free (UAF), and other memory corruption violations (MEMCRP), such as arbitrary memory accesses and dropping of uninitialized memory. Out of the 42 found memory corruption vulnerabilities, 30 are previously-known which deepSURF detects automatically without any human involvement. In addition, deepSURF discovers 12 new memory bugs in seven crates presented in Table III. We have disclosed all the newly-discovered bugs to their affected developers. After our reporting, three bugs have been patched and 11 have been assigned RustSec IDs.
Many of the memory bugs detected by deepSURF involve intricate scenarios and advanced Rust features that pose challenges for automated tools. For example, triggering 21 bugs requires custom function implementations that panic or return unexpected values. Others, such as those in slice-deque [^2] The crate slice-deque has been unmaintained since 2020, but the bugs we found also affect its maintained fork: slice-ring-buffer. and simple-slab, are triggered only by specific API call sequences. These bug-triggering patterns are typically found only in human-written PoCs, yet deepSURF generated harnesses containing them automatically. Section 4 provides more details about these harnesses.
deepSURF misses 26 reported memory bugs across 22 crates from the ERAsan and RustSan datasets. Several of these require language features beyond its current scope, including async programming, multithreading, or multi-step sequences that combine library APIs with internal operations (e.g., calls to std::mem::forget).
RQ1: deepSURF identified 42 memory safety bugs (including 12 previously-unknown) by fuzzing the harnesses that it automatically generated.
RQ2: Comparison with State-of-the-Art Tools. We evaluated existing Rust fuzzing tools on ERAsan’s dataset to compare their bug-finding capabilities with deepSURF. This dataset is a subset of our dataset, including 27 crates with well-documented memory safety issues. Specifically, we ran RUG [12], RPG [64] and RULF [26] to generate harnesses for the libraries in that dataset and fuzzed these targets to compare their findings with deepSURF. For this comparison, the other tools were tested with their original settings, while deepSURF as described in RQ1. We could not test RuMono [69] and FRIES [66], as their code was not publicly available at the time of writing.
As shown in Table IV, deepSURF detected 26 memory safety bugs in ERAsan’s dataset, including 6 previously-unknown. None of the other tools were able to automatically expose any memory corruption bugs in the same dataset. Moreover, deepSURF achieved a high overall URAPI Coverage (see Section 2) of 87.3%, while RUG, RPG, and RULF achieved just 22.5%, 4.1%, and 3% respectively. This significant difference highlights deepSURF’s focus on fuzzing unsafe code reachable through APIs, along with its broad support for Rust data types, enabling it to cover a substantially larger set of potentially vulnerable APIs.
RUG was the best-performing tool among the existing ones, generating harnesses of relatively high quality. While it failed to detect any memory corruption vulnerabilities when tested with its default fuzzing engine (libFuzzer [30]), we observed that it could detect one of the bugs by just switching to AFL++ and potentially seven more after modifying its prompt. In contrast, RULF and RPG exhibit limited argument-type support, preventing them from fully addressing C2 and C3. As a result, their ability to fuzz complex libraries with diverse API argument types is limited.
TABLE IV: Bug-finding capability of deepSURF and comparison of URAPI Coverage and number of compilable generated harnesses with other Rust fuzzing tools. — signifies cases where the tool crashes and ! denotes cases where harness generation did not finish within 24 hours.
@@ 6.2 Evaluation Results
string-interner 0 100% (2) 33.3% (8) 33.3% (1) 33.3% (1)
http 0 86.4% (109) 52.5% (173)
@@
Total 26 (6) 87.3% (1079) 22.5%(636) 4.1% (129) 3% (72)
None of the existing tools fully addresses C4 (support for custom implementations simulating user-defined logic) or C5 (support for generating semantically-related complex API sequences), which are critical for detecting complex memory corruption vulnerabilities. Moreover, all three existing tools failed to properly address C1, since they produced false positives due to crashes unrelated to actual bugs either by directly invoking unsafe code without preserving safety invariants or due to incorrect handling of crashes caused by developer-introduced panics and assertions. Finally, RUG and RPG failed to run on five crates, and RULF on one, due to crashes, timeouts, or errors caused by unimplemented features or implementation errors.